OpenStudy (anonymous):

to solve by completing the square, what value should i add to each side of the equation: x^2+16x=-1 i chose 1...is that correct?

OpenStudy (anonymous):

no... that left side needs to "become" a perfect square trinomial...

OpenStudy (anonymous):

64

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes.. that's what's needed to make the left side a perfect square trinomial... so the right side will be -1 + 64 = ???

OpenStudy (anonymous):

63

OpenStudy (anonymous):

ok... so you have: \(\large x^2+16x+64=63 \) \(\large (x+??)^2=63 \) what goes in place of ???

OpenStudy (anonymous):

8

OpenStudy (anonymous):

\(\large (x+8)^2=63 \) take the square root of both sides and isolate the x...